# Climbing over the potential barrier during inflation via null energy condition violation

### v1

@@ I Introduction
Attributing the observed signal in the PTA band to primordial GWs necessitates physics beyond standard slow-roll inflation, see, e.g., Jiang:2023gfe ; Zhu:2023lbf ; Ye:2023tpz ; Borah:2023sbc ; Choudhury:2023kam ; Ben-Dayan:2023lwd ; Oikonomou:2023qfz ; Datta:2023xpr , for recent studies. There have been numerous studies exploring the generation of a blue-tilted GW spectrum by introducing new physics beyond conventional slow-roll inflation, e.g., Piao:2004tq ; Baldi:2005gk ; Piao:2006jz ; Kobayashi:2010cm ; Kobayashi:2011nu ; Endlich:2012pz ; Cai:2014uka ; Gong:2014qga ; Cannone:2014uqa ; Wang:2014kqa ; Kuroyanagi:2014nba ; Cai:2015yza ; Cai:2016ldn ; Wang:2016tbj ; Fujita:2018ehq ; Kuroyanagi:2020sfw ; Akama:2020jko ; Akama:2023jsb ; Giare:2020plo ; Giare:2022wxq ; Oikonomou:2023bli ; Choudhury:2023kam ; Oikonomou:2024aww . An intriguing direction is to introduce a violation of the null energy condition (NEC) during inflation.
Realizing a stable NEC violation is a challenge due to the presence of ghost or gradient instabilities in the primordial perturbations associated with the NEC violation Rubakov:2014jja ; Libanov:2016kfc ; Kobayashi:2016xpl ; Easson:2011zy ; Qiu:2015nha ; Ijjas:2016tpn ; Ijjas:2016vtq ; Dobre:2017pnt . It is first explicitly demonstrated in the framework of effective field theory that a fully stable444“Fully stable” implies that both ghost and gradient instabilities are eliminated throughout the entire history of the universe. NEC violation can be realized in “beyond Horndeski” theories Cai:2016thi ; Creminelli:2016zwa ; Cai:2017tku ; Cai:2017dyi ; Kolevatov:2017voe , see also Cai:2017dxl ; Cai:2017pga ; Mironov:2018oec ; Qiu:2018nle ; Ye:2019frg ; Ye:2019sth ; Mironov:2019qjt ; Akama:2019qeh ; Mironov:2019mye ; Ilyas:2020qja ; Ilyas:2020zcb ; Zhu:2021whu ; Zhu:2021ggm ; Mironov:2022ffa ; Cai:2022ori ; Panda:2024iqu for later developments. On this basis, it has been demonstrated that NEC violation during inflation can have significant observable effects, including: 1) an enhanced and blue-tilted power spectrum of primordial GWs with distinct features Cai:2020qpu ; Cai:2022nqv , which could be detected by PTA Ye:2023tpz , Advanced LIGO and Advanced Virgo Chen:2024mwg , and Taiji Chen:2024jca , 2) the generation of primordial black holes and scalar-induced GWs Cai:2023uhc , 3) an amplified parity-violation effect in primordial GWs Cai:2022lec ; Zhu:2023lhv .
@@ II The model
In our single scalar field model, the universe originates from a slow-roll inflation. As the inflaton rolls slowly down the potential, it encounters a high potential barrier (see Fig. 2 for a sketch of the potential). Typically, a slowly rolling inflaton cannot climb over this high barrier. However, through the mechanism of NEC violation, the inflaton naturally violates the slow-roll conditions for a certain duration, allowing it to climb over the barrier and continue the slow-roll inflation. Such an evolution will lead to interesting observable effects on the primordial perturbations.
@@
where $\phi$ is a dimensionless scalar field, $X=\nabla_{\mu}\phi\nabla^{\mu}\phi$, the EFT operator $L_{\delta g^{00}R^{(3)}}$ is adopted to thoroughly eliminate the ghost and gradient instabilities of the primordial scalar perturbations, see, e.g., Cai:2017dyi for details. For simplicity, we will focus on the primordial GWs in this paper.
@@
Figure 1: The functions $g_{1}(\phi)$ and $g_{2}(\phi)$, where we have set $q_{1}=10$, $q_{2}=6$, $q_{3}=10$, $\phi_{0}=3.2$, $\phi_{3}=-4.38$, $f_{1}=1$, $f_{2}=40$.
@@
where $\lambda$, $m$, $f_{1,2}$, $q_{1,2,3,4}$ and $\phi_{0,1,2,3,4,5}$ are dimensionless constants. We plot $g_{1}(\phi)$, $g_{2}(\phi)$ and $V(\phi)$ for appropriate values of these parameters in Figs. 1 and 2. As the scalar field rolls along the potential from left to right, the universe sequentially experiences the first stage of slow-roll inflation, NEC violation, and the second stage of slow-roll inflation. These two slow-roll stages are labeled as “inf1” and “inf2” respectively.
The parameters introduced above should satisfy $\phi_{3}<\phi_{2}<0<\phi_{1}<\phi_{0}$. As can be seen in Figs. 1 and 2, when $\phi$ is much smaller than $\phi_{3}$, $g_{1}=1$ and $g_{2}=0$, rendering the potential responsible for the first stage of slow-roll inflation. In this regime, the first term of the potential predominates. Conversely, as $\phi$ greatly exceeds $\phi_{0}$, $g_{1}=1$ and $g_{2}=0$, and the potential is responsible for the second stage of slow-roll inflation, with the third term in the potential becoming dominant. The exponential term in each expression serves to modulate the dominance of various terms across different regions of the potential.
@@
We numerically solve the background evolutions and display the evolutions of $\phi$, $\dot{\phi}$, $H$, $(aH)^{-1}$ and $\epsilon\equiv-\dot{H}/H^{2}$ in Figs. 3, 4 and 5. From these figures, we can see that the universe sequentially experiences the first stage of slow-roll inflation, NEC violation, and the second stage of slow-roll inflation. This background evolution is similar to that in the scenario described in Cai:2020qpu . The main difference from the model in Cai:2020qpu lies in the fact that in this model, the energy scales of the first and second stages of slow-roll inflation are closer to each other.
@@
(a)   $H$
(b)   $(aH)^{-1}$
Figure 4: The evolutions of $H$ (left panel) and the conformal Hubble horizon $(aH)^{-1}$ (right panel).
Figure 5: The evolution of $\epsilon=-\dot{H}/H^{2}$. We have $0<\epsilon\ll 1$ during the first and second stages of slow-roll inflation, and $\epsilon\ll-1$ during the NEC-violating stage.
@@ III The power spectrum of Primordial GWs
(a)  $P_{T}$
(b)  $P_{T}/P_{T,\text{inf2}}$
Figure 6: The power spectra $P_{T}$ of primordial GWs for our model obtained by numerically solving Eq. (10). We have set $q_{1}=10$, $q_{2}=6$, $q_{3}=10$, $q_{4}=4$, $\phi_{0}=3.2$, $\phi_{1}=2$, $\phi_{2}=-4$, $\phi_{3}=-4.38$, $\phi_{4}=20$, $\phi_{5}=29$, $f_{1}=1$, $f_{2}=40$, $\lambda=0.01$, $\sigma=23$, $m=-4.5\times 10^{-6}$, $\phi(t_{\rm ini})=-6$, $\dot{\phi}(t_{\rm ini})=0$ and $t_{\rm ini}=0$.
@@
where $P_{T,inf2}={2H_{inf2}^{2}\over\pi^{2}M_{p}^{2}}$. A detailed expression for $\alpha_{4}$ and $\beta_{4}$ can be found in Appendix A. We plot $P_{T}/P_{T,\text{inf2}}=\left|\alpha_{4}-\beta_{4}\right|^{2}$ in Fig. 7 for different choices of $w_{i}$. We can see from Figs. 6(b) and 7(a) that the approximation works well for our model with appropriate choices of $w_{i}$. In our parameterization, the height difference between the two platforms in the potential can be adjusted by changing $\epsilon_{3}=-\dot{H}/H^{2}={3\over 2}(1+w_{j})$.
Figure 7: The power spectrum of the primordial GWs given by Eq. (22). For the left panel, we set $w_{1}=-1$, $w_{2}=-10$, $w_{3}=1$, and $w_{4}=-1$, providing a close approximation to the result shown in Fig. 6(b). In the right panel, we choose $w_{1}=-1$, $w_{2}=-10$, $w_{3}=2/3$, and $w_{4}=-1$, resembling the scenario discussed in Ref. Cai:2020qpu .
@@ IV The energy density spectrum of GWs
$\Omega_{GW}(\tau_{0})=\frac{k^{2}}{12a_{0}^{2}H^{2}_{0}}P_{T}(k)\left[\frac{3\Omega_{{m}}j_{1}(k\tau_{0})}{k\tau_{0}}\sqrt{1.0+1.36\frac{k}{k_{\text{eq}}}+2.50\left(\frac{k}{k_{\text{eq}}}\right)^{2}}\right]^{2},$ (23)
where $H_{0}=67.8\,{\rm km/s/Mpc}$, $\tau_{0}=1.41\times 10^{4}$ Mpc, $a_{0}=1$, $k_{\text{eq}}=0.073\,\Omega_{\text{m}}h^{2}$ Mpc-1 is the wavenumber corresponding to the mode that entered the horizon at the equality of matter and radiation, $\Omega_{\rm m}=\rho_{m}/\rho_{c}$ is the density fraction of matter today, $\rho_{{c}}=3H^{2}_{0}/\big{(}8\pi G\big{)}$ is the critical energy density, and $j_{1}(x)$ is the spherical Bessel function of the first kind. We plot the energy density spectrum of GWs for our model in Figs. 8(a) and 8(b), using both the numerical results obtained by solving Eq. (10) and the analytical results provided by Eq. (22), respectively. Namely, Figs. 8(a) and 8(b) correspond to the power spectra displayed in Figs. 6(a) and 7(a), respectively.
Figure 8: The energy density spectrum of GWs for our model obtained by using both the numerical results obtained by solving Eq. (10) (left panel) and the analytical results provided by Eq. (22) (right panel).
Comparing Figs. 8(a) and 8(b), we observe a good agreement between the energy density spectra obtained by the two methods. In the second case, where the slow-roll parameters are approximated as constants, some details of the mode evolution are not fully reflected. For observational purposes, the distinctive features may include the blue-tilted nature in the middle-frequency range and the decreasing amplitude of these oscillations at higher frequencies.
@@ V Conclusion
In this paper, we propose a new single-field model where the universe sequentially undergoes a first stage of slow-roll inflation, NEC violation, and a second stage of slow-roll inflation. Within the potential responsible for slow-roll inflation, there exists a high barrier, over which the rolling inflaton climb via the NEC violation mechanism before entering the second inflationary stage. Following the NEC violation, there is a process of energy scale reduction, resulting in the energy scales of the two slow-roll inflation stages being relatively close, contrasting with the model in Cai:2020qpu .
@@ Acknowledgements.
Y. C. is supported in part by the National Natural Science Foundation of China (Grant No. 11905224), the China Postdoctoral Science Foundation (Grant No. 2021M692942) and Zhengzhou University (Grant Nos. 32213984, 35220136). Y.-S. P. is supported by NSFC, No.12075246, National Key Research and Development Program of China, No. 2021YFC2203004, and the Fundamental Research Funds for the Central Universities.

### v2

@@ I Introduction
Attributing the observed signal in the PTA band to primordial GWs necessitates physics beyond standard slow-roll inflation, see, e.g., Jiang:2023gfe ; Zhu:2023lbf ; Ye:2023tpz ; Borah:2023sbc ; Ben-Dayan:2023lwd ; Oikonomou:2023qfz ; Datta:2023xpr ; Datta:2023vbs ; Bhattacharya:2023ysp ; Choudhury:2023hfm ; Choudhury:2023kdb ; Choudhury:2024one ; Sharma:2024whg , for recent studies. This would require the amplitude of the primordial GW power spectrum $P_{T}$ to reach a maximum of $P_{T}\simeq 10^{-3}$ at PTA scales (see, e.g., Ye:2023tpz ), where $P_{T}$ is related to the present GW energy density spectrum $\Omega_{\rm{GW}}$ by $P_{T}\simeq 10^{6}\Omega_{\rm{GW}}$. There have been numerous studies exploring the generation of a blue-tilted GW spectrum by introducing new physics beyond conventional slow-roll inflation, e.g., Piao:2004tq ; Baldi:2005gk ; Piao:2006jz ; Kobayashi:2010cm ; Kobayashi:2011nu ; Dudas:2010gi ; Mourad:2017rrl ; Endlich:2012pz ; Cai:2014uka ; Gong:2014qga ; Cannone:2014uqa ; Wang:2014kqa ; Kuroyanagi:2014nba ; Cai:2015yza ; Cai:2016ldn ; Wang:2016tbj ; Fujita:2018ehq ; Kuroyanagi:2020sfw ; Akama:2020jko ; Akama:2023jsb ; Giare:2020plo ; Datta:2022tab ; Giare:2022wxq ; Oikonomou:2023bli ; Choudhury:2023kam ; Oikonomou:2024aww . An intriguing direction is to introduce a violation of the null energy condition (NEC) during inflation.
Realizing a stable NEC violation is a challenge due to the presence of ghost or gradient instabilities in the primordial perturbations associated with the NEC violation Rubakov:2014jja ; Libanov:2016kfc ; Kobayashi:2016xpl ; Easson:2011zy ; Qiu:2015nha ; Ijjas:2016tpn ; Ijjas:2016vtq ; Dobre:2017pnt ; deCesare:2019pqj ; deCesare:2021wmk . It is first explicitly demonstrated in the framework of effective field theory that a fully stable444“Fully stable” implies that both ghost and gradient instabilities are eliminated throughout the entire history of the universe. NEC violation can be realized in “beyond Horndeski” theories Cai:2016thi ; Creminelli:2016zwa ; Cai:2017tku ; Cai:2017dyi ; Kolevatov:2017voe , see also Cai:2017dxl ; Cai:2017pga ; Mironov:2018oec ; Qiu:2018nle ; Ye:2019frg ; Ye:2019sth ; Mironov:2019qjt ; Akama:2019qeh ; Mironov:2019mye ; Ilyas:2020qja ; Ilyas:2020zcb ; Zhu:2021whu ; Zhu:2021ggm ; Mironov:2022ffa ; Cai:2022ori ; Panda:2024iqu for later developments. On this basis, it has been demonstrated that NEC violation during inflation can have significant observable effects, including: 1) an enhanced and blue-tilted power spectrum of primordial GWs with distinct features Cai:2020qpu ; Cai:2022nqv , which could be detected by PTA Ye:2023tpz , Advanced LIGO and Advanced Virgo Chen:2024mwg , and Taiji Chen:2024jca , 2) the generation of primordial black holes and scalar-induced GWs Cai:2023uhc , 3) an amplified parity-violation effect in primordial GWs Cai:2022lec ; Zhu:2023lhv .
@@ II The model
In our single scalar field model, the universe originates from a slow-roll inflation. As the inflaton rolls slowly down the potential, it encounters a high potential barrier (see Fig. 3 for a sketch of the potential). Typically, a slowly rolling inflaton cannot climb over this high barrier. However, through the mechanism of NEC violation, the inflaton naturally violates the slow-roll conditions for a certain duration, allowing it to climb over the barrier and continue the slow-roll inflation, as illustrated by Fig. 1. Such an evolution will lead to interesting observable effects on the primordial perturbations.
Figure 1: In the first panel, inflaton can climb over the barrier which are not too high. But in the other two panels, inflaton requires the NEC violation to climb over the barrier and continue the slow-roll inflation.
@@
where $\phi$ is a dimensionless scalar field, $X=\nabla_{\mu}\phi\nabla^{\mu}\phi$, the EFT operator $L_{\delta g^{00}R^{(3)}}$ is adopted to thoroughly eliminate the ghost and gradient instabilities of the primordial scalar perturbations, see, e.g., Cai:2017dyi for details. According to Cai:2016thi ; Cai:2017dyi , $L_{\delta g^{00}R^{(3)}}$ does not affect the background equations at all, nor does it contribute to the quadratic action of the tensor sector. It only modifies the squared sound speed of the scalar perturbations. For simplicity, we will focus on the primordial GWs in this paper.
@@
Figure 2: The functions $g_{1}(\phi)$ and $g_{2}(\phi)$, where we have set $q_{1}=10$, $q_{2}=6$, $q_{3}=10$, $\phi_{0}=3.2$, $\phi_{3}=-4.38$, $f_{1}=1$, $f_{2}=40$. A sufficiently large positive $g_{2}$ that satisfies $g_{2}{\dot{\phi}}^{4}>M_{\rm P}^{2}\dot{H}$ can prevent ghost instabilities.
@@
where $\lambda$, $m$, $f_{1,2}$, $q_{1,2,3,4}$ and $\phi_{0,1,2,3,4,5}$ are dimensionless constants. We plot $g_{1}(\phi)$, $g_{2}(\phi)$ and $V(\phi)$ for appropriate values of these parameters in Figs. 2 and 3. As the scalar field rolls along the potential from left to right, the universe sequentially experiences the first stage of slow-roll inflation, NEC violation, and the second stage of slow-roll inflation. These two slow-roll stages are labeled as “inf1” and “inf2” respectively.
The parameters introduced above should satisfy $\phi_{3}<\phi_{2}<0<\phi_{1}<\phi_{0}$. As can be seen in Figs. 2 and 3, when $\phi$ is much smaller than $\phi_{3}$, $g_{1}=1$ and $g_{2}=0$, rendering the potential responsible for the first stage of slow-roll inflation. In this regime, the first term of the potential predominates. Conversely, as $\phi$ greatly exceeds $\phi_{0}$, $g_{1}=1$ and $g_{2}=0$, and the potential is responsible for the second stage of slow-roll inflation, with the third term in the potential becoming dominant. The exponential term in each expression serves to modulate the dominance of various terms across different regions of the potential.
@@
The drop of the potential before the hill is significant for the $\phi$ field to climb to the top of the potential. In our model, the equation of motion for $\phi$ is given by Eq. (4). The coefficient in front of $\ddot{\phi}$ is proportional to $Q_{s}H^{2}/\dot{\phi}^{2}$, which must remain positive throughout to avoid ghost instability. To allow the $\phi$ field to climb to the top of the potential without requiring a non-zero initial kinetic energy, we set the model parameters so that $g_{1}<0$ around the NEC-violating phase. This results in the term $3g_{1}H\dot{\phi}$ acting as an anti-friction force, thereby accelerating the $\phi$ field. Therefore, both the magnitude of $\dot{\phi}$ at the beginning of the climb and the coefficient in front of $\ddot{\phi}$ will influence the $\phi$ field’s ability to climb over the hill. This highlights the importance of the drop of the potential before the hill. Otherwise, if $\dot{\phi}$ is too small when the field starts climbing, we might need to finely tune the model parameters to ensure that the $\phi$ field reaches the top of the potential.
We numerically solve the background evolutions and display the evolutions of $\phi$, $\dot{\phi}$, $H$ and $\epsilon\equiv-\dot{H}/H^{2}$ in Figs. 4 and 5. From these figures, we can see that the universe sequentially experiences the first stage of slow-roll inflation, NEC violation, and the second stage of slow-roll inflation. This background evolution is similar to that in the scenario described in Cai:2020qpu . The main difference from the model in Cai:2020qpu lies in the fact that in this model, the energy scales of the first and second stages of slow-roll inflation are closer to each other.
@@
Figure 5: The evolutions of a hundredfold of $H$ (blue curve) and $\epsilon=-\dot{H}/H^{2}$ (red curve). We have $0<\epsilon\ll 1$ during the first and second stages of slow-roll inflation, and $\epsilon\ll-1$ during the NEC-violating stage. Note that when $\dot{\phi}/H$ reaches its peak, $H$ is still far from its maximum.
@@ III The power spectrum of Primordial GWs
Figure 6: The power spectrum $P_{T}$ of primordial GWs for our model obtained by numerically solving Eq. (10). We have set $q_{1}=10$, $q_{2}=6$, $q_{3}=10$, $q_{4}=4$, $\phi_{0}=3.2$, $\phi_{1}=2$, $\phi_{2}=-4$, $\phi_{3}=-4.38$, $\phi_{4}=20$, $\phi_{5}=29$, $f_{1}=1$, $f_{2}=40$, $\lambda=0.01$, $\sigma=23$, $m=-4.5\times 10^{-6}$, $\phi(t_{\rm ini})=-6$, $\dot{\phi}(t_{\rm ini})=0$ and $t_{\rm ini}=0$.
@@
where $P_{T,inf2}={2H_{inf2}^{2}\over\pi^{2}M_{p}^{2}}$. A detailed expression for $\alpha_{4}$ and $\beta_{4}$ can be found in Appendix A. We plot $P_{T}/P_{T,\text{inf2}}=\left|\alpha_{4}-\beta_{4}\right|^{2}$ in Fig. 7 for different choices of $w_{i}$. We can see from Figs. 6 and 7(a) that the approximation works well for our model with appropriate choices of $w_{i}$. In our parameterization, the height difference between the two platforms in the potential can be adjusted by changing $\epsilon_{3}=-\dot{H}/H^{2}={3\over 2}(1+w_{j})$.
Figure 7: The power spectrum of the primordial GWs given by Eq. (22). For the left panel, we set $w_{1}=-1$, $w_{2}=-10$, $w_{3}=1$, and $w_{4}=-1$, providing a close approximation to the result shown in Fig. 6. In the right panel, we choose $w_{1}=-1$, $w_{2}=-8$, $w_{3}=2/3$, and $w_{4}=-1$, resembling the scenario discussed in Ref. Cai:2020qpu .
@@ IV The energy density spectrum of GWs
$\Omega_{\rm{GW}}(\tau_{0})=\frac{k^{2}}{12a_{0}^{2}H^{2}_{0}}P_{T}(k)\left[\frac{3\Omega_{{m}}j_{1}(k\tau_{0})}{k\tau_{0}}\sqrt{1.0+1.36\frac{k}{k_{\text{eq}}}+2.50\left(\frac{k}{k_{\text{eq}}}\right)^{2}}\right]^{2},$ (23)
where $H_{0}=67.8\,{\rm km/s/Mpc}$, $\tau_{0}=1.41\times 10^{4}$ Mpc, $a_{0}=1$, $k_{\text{eq}}=0.073\,\Omega_{\text{m}}h^{2}$ Mpc-1 is the wavenumber corresponding to the mode that entered the horizon at the equality of matter and radiation, $\Omega_{\rm m}=\rho_{m}/\rho_{c}$ is the density fraction of matter today, $\rho_{{c}}=3H^{2}_{0}/\big{(}8\pi G\big{)}$ is the critical energy density, and $j_{1}(x)$ is the spherical Bessel function of the first kind. We plot the energy density spectrum of GWs for our model in Fig. 8, using the numerical results obtained by solving Eq. (10). Namely, Fig. 8 correspond to the power spectrum displayed in Fig. 6.
Figure 8: The energy density spectrum of GWs for our model. The blue curve is obtained by using the numerical results obtained by solving Eq. (10), while the red curve is based on the power spectrum shown in Fig. 7(b). We have set $P_{T,inf2}=1.82\times 10^{-10}$.
For observational purposes, the distinctive features may include the blue-tilted nature in the middle-frequency range and the decreasing amplitude of these oscillations at higher frequencies. Recently, observational data reported by NANOGrav NANOGrav:2023gor ; NANOGrav:2023hvm ; NANOGrav:2020bcs , EPTA EPTA:2023fyk , PPTA Reardon:2023gzh and CPTA Xu:2023wog is consistent with a stochastic GW background with an amplitude of $\Omega_{\rm{GW}}\simeq 10^{-9}$ and a spectral index of $n_{T}=1.8\pm 0.3$ at the reference frequency $f=1$ $\text{yr}^{-1}$. If the observational data are interpreted as primordial GW background, it corresponds to a strongly blue-tilted spectrum with an amplitude of $P_{T}\simeq 10^{6}\Omega_{\rm{GW}}\simeq 10^{-3}$ at a frequency of $f=1$ $\text{yr}^{-1}$. We have plotted the energy density spectrum in red in Fig. 8 using the analytical formula given in Eq. (22) for a specific set of parameters. Fig. 8 suggests that future PTA observations may have the potential to detect or constrain our model through the specific features of the power spectrum, which requires further investigation.
@@ V Conclusion
In this paper, we propose a new single-field model where the universe sequentially undergoes a first stage of slow-roll inflation, NEC violation, and a second stage of slow-roll inflation. Within the potential responsible for slow-roll inflation, there exists a high barrier, over which the rolling inflaton climb via the NEC violation mechanism [^5] See also Caravano:2024tlp for other mechanism. before entering the second inflationary stage. Following the NEC violation, there is a process of energy scale reduction, resulting in the energy scales of the two slow-roll inflation stages being relatively close, contrasting with the model in Cai:2020qpu .
@@ Acknowledgements.
Y. C. is supported in part by the Natural Science Foundation of Henan Province and Zhengzhou University (Grant Nos. 242300420231, JC23149007, 35220136), the China Postdoctoral Science Foundation (Grant No. 2021M692942). Y.-S. P. is supported by NSFC, No.12075246, National Key Research and Development Program of China, No. 2021YFC2203004, and the Fundamental Research Funds for the Central Universities.